# Exploring the Most Extreme Gamma-Ray Blazars Using Broadband Spectral Energy Distributions

### v1

@@ 4.2 Host galaxy classification
After modelling all the sources in the sample, we find that the best-fit host galaxy model for 59 sources corresponds to the lenticular galaxy template, while for 54 sources, it corresponds to an elliptical galaxy. Among these, 27 sources match an elliptical galaxy of 13 Gyr, 10 sources match an elliptical galaxy of 5 Gyr, and 17 sources match an elliptical galaxy of 2 Gyr. We note that, although the fitting is done using four different host galaxy templates, it is often difficult to distinguish between the three elliptical galaxy templates, as seen in Figure 4. The best-fit host galaxy results given in Table LABEL:table:bestfit correspond to the host galaxy template that yields the lowest $\chi^{2}$ value, but the difference between the different elliptical galaxy templates can be minimal. For the temperature of the host galaxy, we obtain values ranging from $3.2\times 10^{3}$ K to $6.4\times 10^{3}$ K.
Figure 10: CD distribution of the resulting lenticular (blue) and elliptical host galaxies (orange) in the sample.
In our broadband SED modelling, we use templates for elliptical and lenticular galaxies to estimate the thermal emission from the host galaxy. Figure 10 shows that no significant differences are observed between the two types of galaxies. This supports the conclusion that the emission from the host galaxy has a negligible impact on the blazar’s non-thermal emission. Even in galaxies with high star formation rates, the thermal emission is unlikely to significantly contribute to the high-energy regime, particularly for EHSPs. This is consistent with the findings of Scarpa et al. (2000) and Ghisellini et al. (2011), who argue that the thermal emission from the host galaxy is generally not a significant source of seed photons for the external Compton mechanism. While external Compton on star-forming photons could contribute additional components in some extreme cases, like M82 (Ghisellini et al., 2017), we find that this effect does not substantially alter the SEDs of the majority of EHSPs.
@@ 4.3 Energy budget
JetSeT modelling also provides an energy report for the resulting best-fit model, including the magnetic and kinetic energy densities, $U_{B}$ and $U_{e}$, respectively. $U_{B}$ quantifies the energy stored in the magnetic field per unit volume, while $U_{e}$ represents the energy carried by the population of relativistic electrons, often derived from the modelled electron distribution in the jet. Additional outputs from JetSeT include the luminosity of each emission process, i.e., synchrotron and IC, and the jet luminosity associated with radiative mechanisms, electrons, and the magnetic field. The ratio $U_{B}/U_{e}$, representing the energy density of the magnetic field to that of the relativistic electrons, is a critical parameter for understanding the physical conditions and energy balance within blazar jets. This quantity reveals the jet’s magnetisation and the balance between magnetic fields and particle energies in producing the observed radiation. A jet close to equipartition ($U_{B}/U_{e}\sim 1$) is thought to be energetically efficient, as such a configuration minimizes energy losses during the acceleration and transport of particles. In the context of EHSPs, studying $U_{B}/U_{e}$ helps to probe the mechanisms responsible for their extreme synchrotron peak frequencies, such as efficient particle acceleration and strong magnetic fields. Furthermore, deviations from equipartition can indicate variations in jet dynamics, such as regions of high particle dominance during flaring states or magnetic dominance in steady emission. These characteristics are crucial for modelling the jets and understanding the interplay between synchrotron and inverse Compton processes, as well as the overall energy budget of the system. These energy budget results are given in Table LABEL:table:energy-budget. Figure 11 illustrates the relationship between the magnetic and kinetic energy densities resulting from the SED modelling for the 113 sources in the sample with a good fitting result ($\chi^{2}/\text{dof}<1.5$). For comparison, the results for the magnetic and kinetic energy densities from Tavecchio & Ghisellini (2016), Nievas Rosillo et al. (2022), and Zhao et al. (2024) are also shown.
@@
Interestingly, our results suggest a relation between the $U_{B}/U_{e}$ ratio and the synchrotron peak frequency, with the least energetic sources having lower values of $U_{B}/U_{e}$, and the most extreme sources located closer to the line $U_{B}\approx U_{e}$. To verify if there is a dependence of the $U_{B}/U_{e}$ ratio on the synchrotron peak frequency, we divide our sample into three subsamples with different ranges of $\nu_{sync}^{peak}$ values obtained from our SED modelling. The three subsamples have approximately the same number of sources, and their $U_{B}/U_{e}$ distributions are shown in Figure 11. We fit each of the three distributions of the $U_{B}/U_{e}$ ratio to a Gaussian function, obtaining the following results for the mean and standard deviation: for the distribution of sources with $\nu_{sync}^{peak}\leq 6.7\times 10^{16}$ Hz, $\langle\log(U_{B}/U_{e})\rangle=-1.89$, $\sigma=1.30$; for the distribution of sources with $6.7\times 10^{16}$ Hz $<\nu_{sync}^{peak}\leq 4.3\times 10^{17}$ Hz, $\langle\log(U_{B}/U_{e})\rangle=-1.45$, $\sigma=1.08$; and for the distribution of sources with $\nu_{sync}^{peak}>4.3\times 10^{17}$ Hz, $\langle\log(U_{B}/U_{e})\rangle=-0.90$, $\sigma=1.40$. Therefore, we conclude that the most extreme sources are generally closer to equipartition than the sources with lower synchrotron peak frequencies. This is in agreement with the results obtained by Nievas Rosillo et al. (2022), where most of the sources in their sample, which contains 17 EHSP candidates out of the total 22 sources, were close to equipartition.
@@
However, most sources from the samples of Zhao et al. (2024) and Tavecchio & Ghisellini (2016) are far from equipartition, with the magnetic energy density being much smaller than the electron energy density, clustering mainly around the line $U_{B}=10^{-2}\times U_{e}$. The source sample used by Zhao et al. (2024) contains 348 HSP blazars. Their initial sample was selected by collecting all the blazars classified as HBL in the 4FGL catalogue. Afterwards, they modelled the broadband SED and selected only those blazars with a synchrotron peak frequency of $\nu_{\text{sync}}^{\text{peak}}\geq 10^{15}$ Hz. Among these resulting 348 sources, 42 are also included in our sample. On the other hand, Tavecchio & Ghisellini (2016) selected 45 BL Lac objects, 12 of which were detected in the TeV $\gamma$-ray band. Note that sources detected in the TeV band may show biased results since they are typically detected during flaring episodes. These sample selections can help explain the differences in the obtained results, since the source samples from Tavecchio & Ghisellini (2016) and Zhao et al. (2024) also include variable sources. This variability suggests that during certain observations, such sources may be far from equilibrium (e.g., during flares), and in those cases, the electron energy injection could be higher. This would lead to a lower magnetisation, and consequently, to a lower $U_{B}/U_{e}$ ratio. In contrast, the sources in our sample are characterised by low variability across various wavelengths, due to the selection criteria applied, and it is expected that they are closer to equipartition.
@@
Figure 12: CD as a function of the ratio between the magnetic and the kinetic energy density.
@@ 4.4 The blazar sequence and the role of EHSP blazars
However, the large number of EHSPs we identify may challenge previous assumptions about this source type. Instrumental sensitivity limits, biased selection criteria, incomplete data coverage, and theoretical models may not have fully captured the diversity of blazars (e.g., Ghisellini, 2016; Biteau et al., 2020). If EHSPs are more common than previously thought, the sharp dichotomy between low-luminosity, high-peak blazars and their luminous, low-peak counterparts could soften, suggesting a more continuous distribution. This shift challenges the sequence’s traditional boundaries and invites consideration of alternative interpretations. Furthermore, the difficulty of detecting EHSPs at high redshifts due to their lower luminosity introduces a selection bias (e.g., Giommi et al., 2012; Paliya, 2015; Ajello et al., 2022), potentially skewing the observed relationship between luminosity and synchrotron peak frequency.
The trends in CD provide additional context. The prevalence of low CD values among EHSPs suggests that synchrotron/SSC processes dominate their emission, consistent with the absence of strong external photon fields (Paliya et al., 2021). This contrasts with the high CD values typically seen in FSRQs, where external Compton scattering plays a significant role (e.g., Sol & Zech, 2022). Comparing the CD distribution of EHSPs to other blazar types (see Figure 9) highlights the distinct physical environments and emission mechanisms across the blazar population.
EHSPs, with their extreme synchrotron peaks and low CD, may represent either the natural extension of the blazar sequence or a distinct subclass that deviates from its predictions. Some of the most extreme EHSPs in the sample, particularly those with the highest peak frequencies and CD $\gtrsim 1$, challenge the traditional framework (e.g., Tavecchio et al., 2010). If their underlying jet physics or environmental conditions differ significantly, they could necessitate a revised model that better captures the diversity of the blazar population. The detectability predictions for EHSPs with the CTAO provide an opportunity to test these ideas. Observations of EHSPs at higher redshifts or with unexpected luminosity characteristics could further refine the blazar sequence or point toward its limitations.
These results suggest that while the blazar sequence provides a useful framework, it may oversimplify the diversity of blazars. EHSPs challenge the universality of the sequence, pointing to the need for models that accommodate the complexities of their environments, emission mechanisms, and population statistics (e.g., Biteau et al., 2020; Acciari et al., 2020). This perspective enriches the understanding of blazars and provides a foundation for exploring their classification and evolution in greater depth.
@@ 5 Detectability predictions with CTAO
Table 1: Expected detection significance with CTAO and redshift of the 20 sources  of the sample that have an expected detection significance $\geq 3\sigma$ with CTAO in 20 hours of observation.  For the sources with an expected detection significance $\geq 3\sigma$, but $<5\sigma$, the time needed  to obtain a $5\sigma$ detection is also given in the fifth column ($t_{5\sigma}$). The sources are sorted by right ascension  (RA) in increasing order. In the redshift reference column, P21  corresponds to Paliya et al. (2021) and G21 to Goldoni (2021).
4FGL Name CTAO ($\sigma$) Redshift Redshift reference $t_{5\sigma}$ [h]
@@
Characterised by spectra extending to extremely high frequencies, EHSPs are generally regarded as promising VHE emitters (e.g., Foffano et al., 2019; Acciari et al., 2020; Zhu et al., 2021). Since the higher-energy component of the SED peaks in the VHE range for many EHSPs, studying these sources at TeV energies ($\sim 10^{26}$ Hz) is essential to understand the acceleration processes occurring in their jets. However, their low fluxes, relatively steady emission, and EBL attenuation at TeV energies make their detection in the VHE band challenging. As a result, only a few EHSPs have been observed above 100 GeV, mostly by the current generation of IACTs: MAGIC, VERITAS, and H.E.S.S. The upcoming Cherenkov Telescope Array Observatory (CTAO; Cherenkov Telescope Array Consortium et al., 2019) will represent the next generation of IACTs. These telescopes detect the faint flashes of Cherenkov light produced by the interaction of the $\gamma$ rays with the atmosphere and can only operate during the night, resulting in low duty cycles. Consequently, selecting promising targets for VHE detection is important to optimise the observation time of IACTs. Observations by arrays such as LHAASO or SWGO could also complement these observations for the highest-energy $\gamma$-ray emitters.
@@
Using the spectral shape resulting from the SED modelling explained in Section 3, we estimate the expected detection significance of the modelled sources with the future CTAO assuming a total observation time of 20 hours, following the procedure presented in Domínguez et al. (2024a). For this, our best-fit broadband model is extrapolated to TeV energies, incorporating $\gamma$-ray absorption due to EBL interactions using the optical depths from Domínguez et al. (2024b), based on the EBL model by Saldana-Lopez et al. (2021). The significance estimation for each source is done using the CTAO instrument response functions [^5] https://zenodo.org/records/5499839 (IRFs) for the Alpha Configuration with an average azimuth angle. These IRFs, produced for three different zenith angles (20 degrees, 40 degrees, and 60 degrees), are available for both the north and south hemisphere CTAO arrays. The CTAO-North IRFs are used for sources with positive declination, while the CTAO-South IRFs are used for sources with negative declination. The zenith angle at which each source is observed is determined by assuming observations around culmination time, and the IRF configuration corresponding to the closest zenith angle is used for the significance estimation. Note that the energy threshold is taken as the energy of the lowest bin of the differential sensitivity curve, meaning that the threshold is higher for larger zenith angle observations and for sources in the southern hemisphere. This is because the CTAO northern array, equipped with larger telescopes, is optimized for greater sensitivity to lower-energy $\gamma$-rays, whereas the southern array is designed to detect the highest-energy Cherenkov showers.
Following the procedure by Domínguez et al. (2024a), point-like differential sensitivity curves for a total exposure time of 20 hours are generated by scaling the IRFs corresponding to a 5-hour exposure with gammapy (Deil et al., 2017; Acero et al., 2022). From this analysis, we derive the differential flux, as well as the number of excess and background events required to generate a $5\sigma$ signal within each energy bin. This number of excess events is then scaled linearly with the ratio of the differential fluxes, which is the differential flux in each energy bin for the assumed spectral shape of the source divided by the differential flux derived from the sensitivity curve. As a result, the number of excess events for the considered source is obtained. Finally, the expected detection significance for each source is estimated by applying the equation from Li & Ma (1983) to the obtained number of excess and background events, using the WStatCountsStatistic function in gammapy.
@@
Out of the 113 blazars in the sample, 9 sources appear to be potential VHE emitters, with an expected detection significance $\geq 5\sigma$ after 20 hours of exposure with CTAO. There are 11 additional sources with an expected detection significance $\geq 3\sigma$ that could be detected with a longer exposure (see Table 1). According to our SED modelling results, of the 20 sources detectable above $3\sigma$, 12 sources have $\nu_{sync}^{peak}\geq 10^{17}$ Hz. Additionally, all these sources are characterised by very low magnetisation, with magnetic field values ranging from 0.0084 to 0.74 Gauss. However, we observe no clear relation between the $U_{B}/U_{e}$ ratio and their detectability predictions with the future CTAO, adding to the findings of Nievas Rosillo et al. (2022), who identified the two potential VHE emitters in their sample as having the smallest $U_{B}/U_{e}$ ratios.

### v2

@@ 4.2 Host galaxy classification
After modelling all the sources in the sample, we find that the best-fit host galaxy model for 59 sources corresponds to the lenticular galaxy template, while for 54 sources, it corresponds to an elliptical galaxy. Among these, 27 sources match an elliptical galaxy of 13 Gyr, ten sources match an elliptical galaxy of 5 Gyr, and 17 sources match an elliptical galaxy of 2 Gyr. We note that, although the fitting was done using four different host galaxy templates, it is often difficult to distinguish between the three elliptical galaxy templates, as seen in Figure 4. The best-fit host galaxy results given in Table LABEL:table:bestfit correspond to the host galaxy template that yielded the lowest $\chi^{2}$ value, but the difference between the different elliptical galaxy templates can be minimal. For the temperature of the host galaxy, we obtain values ranging from $3.2\times 10^{3}$ K to $6.4\times 10^{3}$ K.
Figure 10: Compton dominance distribution of the resulting lenticular (blue) and elliptical host galaxies (orange) in the sample.
In our broadband SED modelling, we used templates for elliptical and lenticular galaxies to estimate the thermal emission from the host galaxy. Figure 10 shows that no significant differences are observed between the two types of galaxies. This supports the conclusion that the emission from the host galaxy has a negligible impact on the blazar’s non-thermal emission. Even in galaxies with high star formation rates, the thermal emission is unlikely to significantly contribute to the high-energy regime, particularly for EHSPs. This is consistent with the findings of Scarpa et al. (2000) and Ghisellini et al. (2011), who argued that the thermal emission from the host galaxy is generally not a significant source of seed photons for the external Compton mechanism. While external Compton on star-forming photons could contribute additional components in some extreme cases, like M82 (Ghisellini et al., 2017), we find that this effect does not substantially alter the SEDs of the majority of EHSPs.
@@ 4.3 Energy budget
JetSeT modelling also provides an energy report for the resulting best-fit model, including the magnetic and kinetic energy densities, $U_{B}$ and $U_{e}$, respectively. The former quantifies the energy stored in the magnetic field per unit volume, while the latter represents the energy carried by the population of relativistic electrons, often derived from the modelled electron distribution in the jet. Additional outputs from JetSeT include the luminosity of each emission process, i.e. synchrotron and IC, and the jet luminosity associated with radiative mechanisms, electrons, and the magnetic field. The ratio $U_{B}/U_{e}$, representing the energy density of the magnetic field to that of the relativistic electrons, is a critical parameter for understanding the physical conditions and energy balance within blazar jets. This quantity reveals the jet’s magnetisation and the balance between magnetic fields and particle energies in producing the observed radiation. A jet close to equipartition ($U_{B}/U_{e}\sim 1$) is thought to be energetically efficient, as such a configuration minimises energy losses during the acceleration and transport of particles. In the context of EHSPs, studying $U_{B}/U_{e}$ helps to probe the mechanisms responsible for their extreme synchrotron peak frequencies, such as efficient particle acceleration and strong magnetic fields. Furthermore, deviations from equipartition can indicate variations in jet dynamics, such as regions of high particle dominance during flaring states or magnetic dominance in steady emission. These characteristics are crucial for modelling the jets and understanding the interplay between synchrotron and inverse Compton processes, as well as the overall energy budget of the system. These energy budget results are given in Table LABEL:table:energy-budget. Figure 11 illustrates the relationship between the magnetic and kinetic energy densities resulting from the SED modelling for the 113 sources in the sample with a good fitting result ($\chi^{2}/\text{dof}<1.5$). For comparison, the results for the magnetic and kinetic energy densities from Tavecchio & Ghisellini (2016), Nievas Rosillo et al. (2022), and Zhao et al. (2024) are also shown.
@@
Interestingly, our results suggest a relation between the $U_{B}/U_{e}$ ratio and the synchrotron peak frequency, with the least energetic sources having lower values of $U_{B}/U_{e}$, and the most extreme sources located closer to the line $U_{B}\approx U_{e}$. To verify if there is a dependence of the $U_{B}/U_{e}$ ratio on the synchrotron peak frequency, we divided our sample into three subsamples with different ranges of $\nu_{sync}^{peak}$ values obtained from our SED modelling. The three subsamples have approximately the same number of sources, and their $U_{B}/U_{e}$ distributions are shown in Figure 11. We fitted each of the three distributions of the $U_{B}/U_{e}$ ratio to a Gaussian function, obtaining the following results for the mean and standard deviation: for the distribution of sources with $\nu_{sync}^{peak}\leq 6.7\times 10^{16}$ Hz, $\langle\log(U_{B}/U_{e})\rangle=-1.89$, $\sigma=1.30$; for the distribution of sources with $6.7\times 10^{16}$ Hz $<\nu_{sync}^{peak}\leq 4.3\times 10^{17}$ Hz, $\langle\log(U_{B}/U_{e})\rangle=-1.45$, $\sigma=1.08$; and for the distribution of sources with $\nu_{sync}^{peak}>4.3\times 10^{17}$ Hz, $\langle\log(U_{B}/U_{e})\rangle=-0.90$, $\sigma=1.40$. Therefore, we conclude that the most extreme sources are generally closer to equipartition than the sources with lower synchrotron peak frequencies. This is in agreement with the results obtained by Nievas Rosillo et al. (2022), where most of the sources in their sample, which contains 17 EHSP candidates out of the total 22 sources, were close to equipartition.
@@
However, most sources from the samples of Zhao et al. (2024) and Tavecchio & Ghisellini (2016) are far from equipartition, with the magnetic energy density being much smaller than the electron energy density, clustering mainly around the line $U_{B}=10^{-2}\times U_{e}$. The source sample used by Zhao et al. (2024) contains 348 HSP blazars. Their initial sample was selected by collecting all the blazars classified as HBL in the 4FGL catalogue. Afterwards, they modelled the broadband SED and selected only those blazars with a synchrotron peak frequency of $\nu_{\text{sync}}^{\text{peak}}\geq 10^{15}$ Hz. Among these resulting 348 sources, 42 are also included in our sample. On the other hand, Tavecchio & Ghisellini (2016) selected 45 BL Lac objects, 12 of which were detected in the tera-electronvolt $\gamma$-ray band. Note that sources detected in the tera-electronvolt band may show biased results since they are typically detected during flaring episodes. These sample selections can help explain the differences in the obtained results, since the source samples from Tavecchio & Ghisellini (2016) and Zhao et al. (2024) also include variable sources. This variability suggests that during certain observations, such sources may be far from equilibrium (e.g. during flares), and in those cases, the electron energy injection could be higher. This would lead to a lower magnetisation, and consequently, to a lower $U_{B}/U_{e}$ ratio. In contrast, the sources in our sample are characterised by low variability across various wavelengths, due to the selection criteria applied, and it is expected that they are closer to equipartition.
Figure 12 illustrates the correlation between the CD and the magnetic to kinetic energy density ratio, showing that sources closer to equipartition tend to have lower CD values. This trend supports the idea that EHSPs typically lack strong external photon fields, and are dominated by SSC emission rather than external Compton processes. The low Compton dominance in these sources indicates that inverse Compton scattering is relatively inefficient, likely due to both weak external fields and a near-equipartition state within the jet. Notably, the observed trend also suggests that more magnetically dominated jets tend to produce less Compton-dominated emission, highlighting the connection between jet magnetisation and the radiative output of the jet.
To test whether the trends observed with synchrotron peak frequency persist when using a more intrinsic physical parameter, we computed for each source the characteristic Lorentz factor, $\gamma_{3P}$, defined as the value that maximises $\gamma^{3}n(\gamma)$ in the best-fit electron energy distribution. While some correlations, such as with the magnetic field or Compton dominance, are also observed with $\gamma_{3P}$, others, most notably the trend between $U_{B}/U_{e}$ and $\nu_{sync}^{peak}$, are not evident when using $\gamma_{3P}$. This suggests that $\nu_{sync}^{peak}$ reflects a more complex combination of parameters (including $\gamma$, B, and $\delta$) and may capture broader physical conditions in the jet.
@@
Figure 12: Compton dominance as a function of the ratio between the magnetic and the kinetic energy density.
@@ 4.4 The blazar sequence and the role of EHSP blazars
However, the large number of EHSPs we identify may challenge previous assumptions about this source type. Instrumental sensitivity limits, biased selection criteria, incomplete data coverage, and theoretical models may not have fully captured the diversity of blazars (e.g. Ghisellini, 2016; Biteau et al., 2020). If EHSPs are more common than previously thought, the sharp dichotomy between low-luminosity, high-peak blazars and their luminous, low-peak counterparts could soften, suggesting a more continuous distribution. This shift challenges the sequence’s traditional boundaries and invites consideration of alternative interpretations. Furthermore, the difficulty of detecting EHSPs at high redshifts due to their lower luminosity introduces a selection bias (e.g. Giommi et al., 2012; Paliya, 2015; Ajello et al., 2022), potentially skewing the observed relationship between luminosity and synchrotron peak frequency.
The trends in CD provide additional context. The prevalence of low CD values among EHSPs suggests that synchrotron/SSC processes dominate their emission, consistent with the absence of strong external photon fields (Paliya et al., 2021). This contrasts with the high CD values typically seen in FSRQs, where external Compton scattering plays a significant role (e.g. Sol & Zech, 2022). Comparing the CD distribution of EHSPs to other blazar types (see Figure 9) highlights the distinct physical environments and emission mechanisms across the blazar population.
With their extreme synchrotron peaks and low CD, EHSPs may represent either the natural extension of the blazar sequence or a distinct subclass that deviates from its predictions. Some of the most extreme EHSPs in the sample, particularly those with the highest peak frequencies and CD 1$\gtrsim$ 1≳ 1, challenge the traditional framework (e.g. Tavecchio et al., 2010). If their underlying jet physics or environmental conditions differ significantly, they could necessitate a revised model that better captures the diversity of the blazar population. The detectability predictions for EHSPs with the CTAO provide an opportunity to test these ideas. Observations of EHSPs at higher redshifts or with unexpected luminosity characteristics could further refine the blazar sequence or point towards its limitations.
These results suggest that while the blazar sequence provides a useful framework, it may oversimplify the diversity of blazars. Blazars of the EHSP class challenge the universality of the sequence, pointing to the need for models that accommodate the complexities of their environments, emission mechanisms, and population statistics (e.g. Biteau et al., 2020; Acciari et al., 2020). This perspective enriches the understanding of blazars and provides a foundation for exploring their classification and evolution in greater depth.
@@ 5 Detectability predictions with CTAO
Table 2: Expected detection significance with CTAO for sources with $\geq 3\sigma$ in 20 hours.
4FGL Name
CTAO
($\sigma$)
Redshift
Redshift
reference
$t_{5\sigma}$ [h]
@@
 [^6] The table gives the expected detection significance with CTAO and the redshift of the 20 sources of the sample that have an expected detection significance $\geq 3\sigma$ with CTAO in 20 hours of observation. For the sources with an expected detection significance $\geq 3\sigma$, but $<5\sigma$, the time needed to obtain a $5\sigma$ detection is also given in the fifth column ($t_{5\sigma}$). The sources are sorted by right ascension (RA) in increasing order. In the redshift reference column, P21 corresponds to Paliya et al. (2021) and G21 to Goldoni (2021).
@@
Characterised by spectra extending to extremely high frequencies, EHSPs are generally regarded as promising VHE emitters (e.g. Foffano et al., 2019; Acciari et al., 2020; Zhu et al., 2021). Since the higher-energy component of the SED peaks in the VHE range for many EHSPs, studying these sources at tera-electronvolt energies ($\sim 10^{26}$ Hz) is essential to understand the acceleration processes occurring in their jets. However, their low fluxes, relatively steady emission, and EBL attenuation at tera-electronvolt energies make their detection in the VHE band challenging. As a result, only a few EHSPs have been observed above 100 GeV, mostly by the current generation of IACTs: MAGIC, VERITAS, and H.E.S.S. The upcoming Cherenkov Telescope Array Observatory (CTAO; Cherenkov Telescope Array Consortium et al., 2019) will represent the next generation of IACTs. These telescopes detect the faint flashes of Cherenkov light produced by the interaction of the $\gamma$ rays with the atmosphere and can only operate during the night, resulting in low duty cycles. Consequently, selecting promising targets for VHE detection is important to optimise the observation time of IACTs. Observations by arrays such as LHAASO or SWGO could also complement these observations for the highest-energy $\gamma$-ray emitters.
@@
Using the spectral shape resulting from the SED modelling explained in Section 3, we estimated the expected detection significance of the modelled sources with the future CTAO assuming a total observation time of 20 hours, following the procedure presented in Dom´ınguez et al. (2024a). For this, our best-fit broadband model was extrapolated to tera-electronvolt energies, incorporating $\gamma$-ray absorption due to EBL interactions using the optical depths from Dom´ınguez et al. (2024b), based on the EBL model by Saldana-Lopez et al. (2021). The significance estimation for each source was done using the CTAO instrument response functions [^7] https://zenodo.org/records/5499839 (IRFs) for the ‘Alpha Configuration’ with an average azimuth angle. These IRFs, produced for three different zenith angles (20 degrees, 40 degrees, and 60 degrees), are available for both the north and south hemisphere CTAO arrays. The CTAO-North IRFs were used for sources with positive declination, while the CTAO-South IRFs were used for sources with negative declination. The zenith angle at which each source is observed was determined by assuming observations around culmination time, and the IRF configuration corresponding to the closest zenith angle was used for the significance estimation. Note that the energy threshold is taken as the energy of the lowest bin of the differential sensitivity curve, meaning that the threshold is higher for larger zenith angle observations and for sources in the southern hemisphere. This is because the CTAO northern array, equipped with larger telescopes, is optimised for greater sensitivity to lower-energy $\gamma$-rays, whereas the southern array is designed to detect the highest-energy Cherenkov showers. Following the procedure by Dom´ınguez et al. (2024a), point-like differential sensitivity curves for a total exposure time of 20 hours were generated by scaling the IRFs corresponding to a 5-hour exposure with gammapy (Deil et al., 2017; Acero et al., 2022). From this analysis, we derived the differential flux, as well as the number of excess and background events required to generate a $5\sigma$ signal within each energy bin. This number of excess events was then scaled linearly with the ratio of the differential fluxes, which is the differential flux in each energy bin for the assumed spectral shape of the source divided by the differential flux derived from the sensitivity curve. As a result, the number of excess events for the considered source was obtained. Finally, the expected detection significance for each source was estimated by applying the equation from Li & Ma (1983) to the obtained number of excess and background events, using the WStatCountsStatistic function in gammapy.
@@
Out of the 113 blazars in the sample, nine sources appear to be potential VHE emitters, with an expected detection significance $\geq 5\sigma$ after 20 hours of exposure with CTAO. There are 11 additional sources with an expected detection significance $\geq 3\sigma$ that could be detected with a longer exposure (see Table 2). According to our SED modelling results, of the 20 sources detectable above $3\sigma$, 12 sources have $\nu_{sync}^{peak}\geq 10^{17}$ Hz. Additionally, all these sources are characterised by very low magnetisation, with magnetic field values ranging from 0.0084 to 0.74 Gauss. However, we observe no clear relation between the $U_{B}/U_{e}$ ratio and their detectability predictions with the future CTAO, adding to the findings of Nievas Rosillo et al. (2022), who identified the two potential VHE emitters in their sample as having the smallest $U_{B}/U_{e}$ ratios.